### v1

# Further author information: (Send correspondence to C.G.A.)

Abstract
The family of multi-plane phase retrieval sensors, such as the curvature and nonlinear curvature wavefront sensors (WFS), contain tip/tilt information embedded in their signals. We have built a nonlinear curvature WFS to study different wavefront reconstruction methods and test the ability to extract tip/tilt information. Using reliable and fast centroiding algorithms, combined with knowledge of the measured -distance to each measurement plane, we demonstrate that image jitter may be sensed and compensated for using a fast steering mirror and the WFS alone, i.e. without the need for peripheral components such as quad-cells or access to a separate scientific imaging channel. This approach, which is both precise and accurate, corroborates previous numerical simulations and is expected to improve the overall reconstruction accuracy of multi-plane phase retrieval sensors including higher order spatial modes.
keywords:
wavefront sensing, adaptive optics, wavefront reconstruction algorithms1 INTRODUCTION
Atmospheric turbulence contains most of its power in the lowest spatial frequencies, manifesting as tip and tilt aberrations when light is collected by a ground-based telescope. Left uncorrected, these low-order modes spatially smear or blur images that would otherwise be diffraction-limited; in the case of spectroscopy, significant signal may be lost when the target drifts beyond the limits of a slit or fiber [1, 2]. Residual tip/tilt errors can also degrade the closed-loop reconstruction accuracy of higher-order aberrations of Adaptive Optics (AO) systems[3]. With multi-plane diffractive WFS, such as the nonlinear Curvature Wavefront Sensor (nlCWFS), Abbott et al. 2025 have shown that accurate recovery of high-order spatial modes depends critically on precise tip/tilt sensing and jitter control [4].
The nlCWFS offers excellent sensitivity and resilience to scintillation compared with other WFS such as the Shack-Hartmann Wavefront Sensor (SHWFS) [5, 6, 7, 8, 9]. Typically four measurement planes are employed, each offset from the pupil along the optical axis, and a phase retrieval algorithm is used to reconstruct the wavefront phase and amplitude. Owing to the adaptable placement of -distances at each measurement plane, this architecture provides design flexibility for challenging applications with both monochromatic and broadband light [6, 10, 11, 12].
Jitter sensing can be implemented in the nlCWFS using geometric optics and/or wave optics, enabling tip/tilt correction without ancillary devices and allowing more light to reach scientific instruments[13]. By analyzing spatial offsets in each measurement plane, tip/tilt modes may be traced along the optical path, enabling extraction of their direction and magnitude. A multi-plane configuration provides redundancy and cross-validation as individual planes have tip/tilt information encoded in the image. Small errors in tip/tilt estimation can further cascade into large uncertainties in reconstruction of the overall wavefront, so establishing a reliable and accurate method for extracting tip/tilt is of high priority. Precise jitter control also leads to improved phase unwrapping performance [14].
Abbott et al. (2025) [4] performed simulations evaluating the precision, accuracy, and latency of multiple centroiding methods for the nlCWFS under a range of conditions. By examining residual tip/tilt, they identified two prospective approaches for centroid estimation: a geometric intensity Weighted Average (WA) method and physical-optics-based Speckle Pattern Method (SPM). While SPM proved to be most accurate, achieving better than accuracy (more than three times better than WA for an aberrated beam), it was also significantly more expensive computationally, more than 500 slower. Based on this trade-off, Abbott et al. 2025 recommended the use of fast, empirical algorithms such as WA for real-time applications. The choice of centroiding method thus involves a balance between speed and accuracy, and must ultimately be validated through laboratory experimentation. The objective of this work is to confirm the simulation results reported by Abbott et al. 2025 and demonstrate that an empirical method like WA can indeed achieve diffraction-limited performance as claimed.
To evaluate the simulation results, we have developed a laboratory experiment to compare centroiding precision and accuracy with simulations. Section 2 describes the optical layout, experimental setup, and algorithms used to test tip/tilt analysis and jitter control. Section 3 presents the outcomes of these tests using the intensity weighted average (WA) method and assessing phase reconstruction quality. Finally, Section 4 provides conclusions and a discussion of future directions.
2 EXPERIMENTS AND METHODS
We aim to substantiate the results from Abbott et al, 2025[4] by performing tip/tilt measurements in a lab setting using real hardware in the presence of noise. Experiments conducted involved quantifying centroiding and tip/tilt extraction for unaberrated and aberrated beams. Known amounts of tip/tilt were introduced using a fast steering mirror (FSM) to provide truth measurements for which to compare results.
2.1 Experimental Design
A diagram of the experimental setup is illustrated in Figure 1. Collimated light is injected into the system from a single-mode fiber (SMF) illuminated by a HeNe laser operating at nm. The beam diameter and pupil are defined by a 1 mm pinhole. An aberrating waveplate can optionally be inserted upstream of the pupil to introduce aberrations and scintillation effects. The beam is redirected by a FSM, which is used either to correct tip/tilt offsets for control purposes or introduce known tip/tilt signals for calibration. An optional deformable mirror (DM) is included in the path for closed-loop experiments.111Results for closed-loop control will be reported in a follow-on article. After the DM, the beam passes through a 2:1 beam reducer, decreasing its diameter to 0.5 mm to induce diffraction over short distances. At this scale, the -distances for the four nlCWFS imaging planes are set to cm (near/inner planes) and cm (far/outer planes) as measured from the pupil[15]. Images are recorded using an Allied Vision camera with a plate scale of m. Table 1 lists components in the experiment. Images of key hardware devices are shown in Figure 2.
| Component | Vendor | Model | Notes |
| HeNe Laser | Melles Griot | — | nm |
| FSM | Optics in Motion | OIM5001 | range |
| DM | Boston Micromachines | Multi-C-DM | 140 actuators |
| CCD | Alvium | 1800 U-319m | pixels |
2.2 Centroiding and Calibration
Reliable tip/tilt extraction depends on accurate centroid estimation and image centering. The primary centroiding method used in this study is the intensity weighted-average (WA) method, selected for its balance of speed and accuracy. As reported in Abbott et al. (2025) [4], WA achieved approximately accuracy in 5.7 ms for a array. The centroid location for each image plane (–) is computed as follows:
where the summation runs over all pixel coordinates . Although efficient, WA is known to be sensitive to background noise and intensity outliers. To evaluate its robustness under realistic conditions, our experiments include both unaberrated and distorted beams.
For tip/tilt calibration, a two-step Hough–Canny approach is used to determine the true center of an unaberrated beam in each imaging plane. First, a Canny edge detector identifies the beam’s boundary [16]. Then, a Circular Hough Transform (CHT) fits a circle to the detected edge and extracts the circle’s center as the reference centroid [17]. This method assumes the calibration beam is symmetric and circular, a condition verified in Figure 3.
3 RESULTS
3.1 Tip/Tilt Retrieval
Known amounts of tip and tilt were injected into the optical system using the FSM. Three test cases were evaluated, with offsets of , , and in tip, and corresponding offsets of , , and in tilt. These values were selected randomly to evaluate the accuracy of the retrieval procedure across a moderate dynamic range, while remaining within the operational limits of the FSM and camera. Figure 3 shows the measurement planes and corresponding WA centroid positions for the zero tip/tilt case and the offset condition. The accuracy of the retrieved values is shown in Figure 4, which plots the measured versus injected phase along with the residuals. For an otherwise unaberrated beam, the WA algorithm, when averaging the results from the outer planes, is able to estimate the tip and tilt to within on average across the three cases. In contrast, the inner planes exhibit a wider spread in their estimates () and are less reliable for tip/tilt retrieval.
Several sources of statistical and systematic errors may be considered as contributing to uncertainties in measured tip/tilt. One potential contributor is the FSM’s intrinsic accuracy, which is specified as approximately 2% of its full range (equivalent to ). Meanwhile, the device’s angular resolution is much finer, around rad or . The FSM is precise and its accuracy is diffraction-limited over small movements. Given the several angular deviations introduced in our experiment, FSM electro-mechanical functionality is therefore not a dominant contributor to overall performance.
Another source of uncertainty is evident in the reduced accuracy of the inner-plane measurements compared with outer-plane measurements. The subtended angle, , for a region of pixels at a distance from the pupil is
where is the pixel plate-scale. Since the inner-planes are positioned closer to the pupil ( per pixel) than the outer-planes ( per pixel), small angular shifts are more difficult to measure, making near-field tip/tilt estimates sensitive to WA centroiding accuracy (statistical) and optical mis-alignments (systematic). These results support the use of outer-plane data (alone) for tip/tilt estimation due to their finer angular resolution and greater sensitivity to low spatial frequency modes.
3.2 Tip/Tilt Compensation and Reconstruction
To evaluate the impact of tip/tilt detection and correction on wavefront reconstruction, we emulate closed-loop operation using image-domain adjustments. Based on the calculated angular offsets from the outer measurement planes, a corrective pixel shift is applied to each image using Equation 2 by solving for . After applying the correction, phase reconstructions are performed using a modified Gerchberg-Saxton (GS) algorithm with up to five iterations. See Letchev et al. 2024 [9] for a more in-depth discussion of the reconstruction algorithm.
3.2.1 Reconstruction Without Aberrations
To establish a baseline, the tip/tilt correction procedure was first applied to the non-aberrated datasets. Figure 5 shows the phase reconstructions before (top row) and after (bottom row) applying the pixel-space correction derived from the outer-plane estimates. As expected, in the absence of additional aberrations, the correction yields a significantly flattened wavefront, confirming that the tip/tilt estimation and compensation method is functioning as intended.
3.2.2 Reconstruction with Aberrations
To evaluate the effectiveness of tip/tilt correction under more realistic conditions, an aberrator plate was placed near the pupil plane (just upstream of the “telescope”) to induce phase distortions while minimizing scintillation. Figure 6 shows the resulting measurement planes for three different aberration cases. Unlike the known tip/tilt trials, no additional mis-pointing was injected; misalignments arise purely from the aberrated wavefront. A WA centroiding analysis was applied to each spot image, and the resulting centroids are overlaid as red circles in the figure.
The corresponding phase reconstructions, before and after tip/tilt correction, are shown in Figure 7. Improvements are evident in the corrected results, particularly for Aberrations 1 and 3, where the reconstructions no longer exhibit branch cuts and the wavefronts appear significantly smoother. These changes suggest that the tip/tilt estimation and correction procedure is performing effectively. This conclusion is further supported by Figure 8, which shows the root mean square (RMS) wavefront reconstruction error as a function of applied tip/tilt offset. For each aberration, small errors in tip/tilt estimation can lead to large reconstruction inaccuracies. To achieve diffraction-limited performance, the required tip/tilt accuracy is for Aberration 2, and for Aberrations 1 and 3.
Based on prior simulations[4], the WA method is expected to achieve an accuracy of roughly , which is comparable to the requirements estimated from lab data using a single measurement. In practice, integrating WA into a closed-loop AO system that drives tip/tilt modes to null provides a path forward for ensuring reliable wavefront reconstructions. In other words, as tip/tilt correction improves, wavefront reconstructions improve, making the measurement plane images more symmetric, which in turn allows WA to achieve higher precision (down to in the unaberrated case) — establishing a positive feedback loop for the nlCWFS.
4 Summary and Concluding Remarks
This work shows that tip/tilt can be extracted concurrently with high-order phase retrieval using only the internal optics of the nlCWFS, without relying on quad-cells or additional cameras. Building on previous numerical work, which showed tip/tilt reconstruction accuracy of , we have validated theoretical approaches in a laboratory setting, achieving comparable results of for an unaberrated beam and better than in the presence of aberrations. The experiments corroborate that the outer planes provide the best estimate of tip/tilt due to the longer lever arm. Across different experiments, the simple WA centroiding method proved to be an accurate and reliable estimator for detecting tip/tilt offsets.
These results support a vision of a closed-loop AO system in which tip/tilt compensation is integrated directly into the nlCWFS reconstruction pipeline. Such a feedback loop would iteratively improve both wavefront correction and centroid estimation: tip/tilt correction by the FSM flattens the wavefront, leading to more accurate reconstructions, which in turn enhance centroid estimates and further reduce pointing jitter. Future work will focus on implementing this feedback into a real-time FSM and deformable mirror correction system.
Data Availability
The data used to generate figures for this article are available upon request from the authors.
Disclosures
The authors declare there are no financial interests, commercial affiliations, or other potential conflicts of interest that have influenced the objectivity of this research or the writing of this paper.
Acknowledgements.
This research was supported in part by the Air Force Office of Scientific Research (AFOSR) grant number FA9550-22-1-0435 and the Joint Directed Energy Transfer Office (JDETO).References
- [1] Bechter, A. J., Crass, J., Tesch, J., Crepp, J. R., and Bechter, E. B., “Characterization of Single-mode Fiber Coupling at the Large Binocular Telescope,” PASP 132, 015001 (Jan. 2020).
- [2] Crepp, J. R., “Improving planet-finding spectrometers,” Science 346, 809–810 (Nov. 2014).
- [3] Tyson, R. K. and Frazier, B. W., [Principles of Adaptive Optics ], CRC Press, 5th ed. (2022).
- [4] Abbott, C. G., Crepp, J. R., Letchev, S. O., and Smith, C. M., “Centroiding and extraction of tip/tilt information from nonlinear curvature wavefront sensor measurements,” Journal of Astronomical Telescopes, Instruments, and Systems 11(1), 019001 (2025).
- [5] Guyon, O., “High sensitivity wavefront sensing with a nonlinear curvature wavefront sensor,” Publications of the Astronomical Society of the Pacific 122(887), 49–62 (2010).
- [6] Mateen, M., Guyon, O., Sasián, J., Garrel, V., and Hart, M., “A non-linear curvature wavefront sensor reconstruction speed and the broadband design,” in [Astronomical Adaptive Optics Systems and Applications IV ], Tyson, R. K. and Hart, M., eds., 8149, 814909, International Society for Optics and Photonics, SPIE (2011).
- [7] Crass, J., The Adaptive Optics Lucky Imager: combining adaptive optics and lucky imaging, PhD thesis, University of Cambridge (07 2014).
- [8] Crepp, J. R., Letchev, S. O., Potier, S. J., Follansbee, J. H., and Tusay, N. T., “Measuring phase errors in the presence of scintillation,” Opt. Express 28, 37721–37733 (12 2020).
- [9] Letchev, S., Advancing the Technology Readiness of the Nonlinear Curvature Wavefront Sensor through Modeling and Laboratory Investigation, PhD thesis, University of Notre Dame (5 2024).
- [10] Letchev, S., Crass, J., Crepp, J. R., and Potier, S., “Spatial frequency response and sensitivity of the nonlinear curvature wavefront sensor,” in [Adaptive Optics Systems VIII ], Schreiber, L., Schmidt, D., and Vernet, E., eds., Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series 12185, 121858H (Aug. 2022).
- [11] Potier, S., Crepp, J., and Letchev, S., “Developing an error budget for the nonlinear curvature wavefront sensor,” Journal of Astronomical Telescopes, Instruments, and Systems 9, 049004 (Oct. 2023).
- [12] Ahn, K., Guyon, O., Lozi, J., Vievard, S., Deo, V., Lallement, M., and Bragg, J. C., “A non-linear curvature wavefront sensor for the Subaru telescope’s AO3k system,” in [Techniques and Instrumentation for Detection of Exoplanets XI ], Ruane, G. J., ed., 12680, 126800B, International Society for Optics and Photonics, SPIE (2023).
- [13] Abbott, C. G., Crepp, J. R., Letchev, S. O., and Smith, C. M., “Centroiding and extraction of tip/tilt information from nonlinear curvature wavefront sensor measurements,” in [Adaptive Optics Systems IX ], Jackson, K. J., Schmidt, D., and Vernet, E., eds., 13097, 130973X, International Society for Optics and Photonics, SPIE (2024).
- [14] Huerta, D. A., Crepp, J. R., Abbott, C. G., and Joseph, B., “Comparison of Phase-Unwrapping Methods for Adaptive Optics Wavefront Sensing,” SPIE (2025).
- [15] Letchev, S., Crass, J., and Crepp, J., “Assessing phase reconstruction accuracy for different nonlinear curvature wavefront sensor configurations,” Journal of Astronomical Telescopes, Instruments, and Systems 9 (10 2023).
- [16] Canny, J., “A computational approach to edge detection,” IEEE Transactions on Pattern Analysis and Machine Intelligence PAMI-8(6), 679–698 (1986).
- [17] Duda, R. O. and Hart, P. E., “Use of the hough transformation to detect lines and curves in pictures,” Communications of the ACM 15(1), 11–15 (1972).
Caleb G. Abbott is a postdoctoral scholar at the University of Notre Dame. His research has focused on astronomical instrumentation, specifically the advancement of AO technologies. He previously received his PhD degree in astronomy and master’s degree in physics from Georgia State University (2022 and 2019 respectively) and bachelor’s degrees in astronomy and physics from the University of Michigan (2014).
Justin R. Crepp is an experimental astrophysicist at the University of Notre Dame. He designs and builds instruments for ground-based telescopes. His research focuses on developing technologies related to adaptive optics. Prior to teaching at Notre Dame, he was a postdoctoral scholar at the California Institute of Technology (2008 to 2012). He received his PhD in astronomy from the University of Florida in 2008 and a bachelor’s degree in physics from the Pennsylvania State University in 2003.
Brian Sands is a senior research engineer with the Engineering and Design Core Facility at the University of Notre Dame. His current areas of expertise are in control system design and integration for ground-based astronomical instruments, and development of hardware-based real-time control systems. Prior to his position at Notre Dame, he was a research scientist at the Air Force Research Laboratory (2005-2017) focusing on gas discharge physics and atmospheric-pressure plasmas. He received his master’s degree in Physics from Miami University in 2005 and a bachelor’s degree in Computational Physics from Pennsylvania State University in 2002.